OpenStudy (anonymous):

Which line has an equation of y=-5x+4 in slope-intercept form? a line passing through the points (1, 9) and (3, 19) a line passing through the points (2, –14) and (4, –24) a line passing through the points (1, 1) and (3, 11) a line passing through the points (2, –6) and (4, –16)

OpenStudy (anonymous):

2nd Option

OpenStudy (akashdeepdeb):

Do you know the slope-intercept form of a line? y = mx + b [Where m is slope and b is y-intercept]

OpenStudy (anonymous):

yes

OpenStudy (akashdeepdeb):

So here in this question has a slope of -5.

OpenStudy (anonymous):

\[\frac{ -14-(-24) }{ 2-4 }\] \[\frac{ -14+24 }{ 2-4 }\] \[\frac{ 10 }{ -2}\] That's the only option Where slope = -5

OpenStudy (akashdeepdeb):

Find the slope of all other lines and see which one has a slope of -5

OpenStudy (akashdeepdeb):

All of them do. :D :p

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

lol

OpenStudy (akashdeepdeb):

@FilthyMcNasty Even the last one does.

OpenStudy (anonymous):

Oh yeahh??

OpenStudy (anonymous):

that kinda confusing so wait the answer is B

OpenStudy (akashdeepdeb):

No it is either B or D. So how do you think we'd decide? :D

OpenStudy (akashdeepdeb):

@andieboo1995 ?

OpenStudy (akashdeepdeb):

@andieboo1995 Did you understand why the answer has to B or D?

OpenStudy (anonymous):

Answer is D.. srry...

OpenStudy (anonymous):

Instead of x,y just use any coordinate & see which 1 is TRUE

OpenStudy (akashdeepdeb):

@andieboo1995 ? Got it then? :)

OpenStudy (anonymous):

y=-5x+4 a line passing through the points (2, –6) and (4, –16) -6= -5(2) +4 -6= -10+4 -6=-6

OpenStudy (anonymous):

ohhhh okay

OpenStudy (anonymous):

lol i didnt get it at first @AkashdeepDeb